OpenStudy (lapshaman):

use the given values to find an equation of the form f(x) ax^2+bx+c f(1)=4, f(2)=12, f(4)=46 I have a math test tomorrow, so some help on this would be great!

OpenStudy (zale101):

I will do the first one f(x)=ax^2+bx+c f(1)=4 means f(x)=4 and the x value is 1. So, replace f(x) with 4 and change the x values to 1 f(1)=4=a(1)^2+b(1)+c 4=a+b+c

OpenStudy (lapshaman):

you need to substitute the numbers for the x's

OpenStudy (lapshaman):

it should be possible............

OpenStudy (yoloshroom):

Ah ok , zale's got this ^_^ he'll clarify. (he said the prev answer is wrong zale.) i think you forgot to say ^2 so it's 4=(1)a^2+b(1)+c(1)

OpenStudy (yoloshroom):

oh no.. hmm, zale should be right... it's a(1)^2 so it's 4=a+b+c...

OpenStudy (lapshaman):

the answer should be a trinomial......

OpenStudy (yoloshroom):

hmm @Zale101

OpenStudy (lapshaman):

my teacher says these are monster problems........

OpenStudy (zale101):

Lapshaman, i did the first equation for you. In the beginning, i made it clear that i left the rest for you to continue.

OpenStudy (zale101):

So please, participate and engage in your question so you can make your learning process easier to grasp

OpenStudy (lapshaman):

I don't think you are doing it right tho, all three need to be compared at the same time, something like: a+b+c=4 4a+2b+c=12 16a+4b+c=46

OpenStudy (zale101):

Correct. You can either use the elimination or the substitution method.

OpenStudy (lapshaman):

how?

OpenStudy (zale101):

Using the elimination method equation 1: a+b+c=4 equation 2: 4a+2b+c=12 equation 3: 16a+4b+c=46 subtract eq 1 and 2 together and subtract 1 and 3 together After doing that', you add both equation together (#1-#2)+(#1-#3)

OpenStudy (lapshaman):

oh, I get it!!

OpenStudy (lapshaman):

that gives me a bunch of negatives.....

OpenStudy (zale101):